Prev Next

SW&IT Meeting Minutes 2018-08-20

Attendees: David, Igor, Kellee, Matthieu, Petr, Yang

Outlook

  • OVMS racktangle
  • AGW4 @ summit
  • LBC dataprobe f/w update (spares)
<!-- for later
* Soul jumbo network
* lbc switches
-->

News from management

  • xxx

Previous Action Items

  • Vsphere (DE)
  • IT email format (DE)
  • LBTI/jumbo switch (DE)
  • power plan update draft (DE)

Open topics

Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.

  • trivial passwords (ovms-uei, others?) (SH when he comes back)
  • wiki software cleanup (MB, mainly due to his own contribution: week reports, etc. - but others too)

This week in github

Team reports

YZ

  • General safety training with David Carroll.
  • Laser safety training with Gustavo.
  • Completed M3 mirror actuator profiling update with Michael Gusick and Michael Gardiner. Updated TCS copy, but yet to push to Github (no write access permission)
  • Finished Python Collimation and plotting tools (with more enhancements to work on).
  • Work on some Sphinx documents examples.
  • Organized current pointing model material (latest Tpoint software, ptconv source code, as well Doug's original scripts), and stored them for now on SVN (under tools/trunk/pointing).
  • Working on a wiki pointing documentation on existing tools and scripts.
  • Continue to study the AO supervisor code and the IT issue #7261 (memory problem with WfsArbitrator).

MB

  • management OC, (no WPC)
  • s/w candidate interview
  • shark-nir follow up with MdP
  • s/w engineer interview
  • pp2018 completion, pp2019 planning
  • power/cooling shutdown activities
  • Argos/Martin followup
  • Laser safety training
  • git migration
next:
  • pp2018 completion, pp2019 planning
  • IT to github issues
  • TCS release planning meet with PK on Aug. 20
  • LBC skytech upgrade meeting with Fernando, touch base on SHARK-VIS

KS

Last week: vacation

This week:
  • OVMS UEI has lost the application files. Is it just a matter of replacing them? What happened? John reseated cards after it wasn't happy with card 8. It's ready to check out.
  • OAC issues? Haven't heard from Dan but MB suggested there were some problems.
  • LBC for restart - put in change to close PDU sockets all the time.

DE

Last week:
  • Surprise power cut at summit
    • Nothing seems to have suffered greatly
    • Shutdown script has some issues
      • Credential problems
        • Installed keys for VM servers
      • Hardware not included in script
        • Synology (ready to go, add to script?)
        • Various instrument computers (Should they be included, or do they need human checks to protect instruments?)
  • Power and VMWare discussions ticking over, deprioritized in favor of summit work
  • AGw computer parts spares in stock
This week:
  • Sort out credentials for shutdown script
  • Find responsible parties and work out the best course of action for instrument computers
  • VMWare licensing
  • Firewall holes for Flightaware computer on summit
  • 4D computer turn on? (Leroy)

IS

  • This week:
    • PIT
      • Mostly General PIT work this week, in particular XSLT PDF work as we are adding a new more concise table of observations and migrating some of the italian fixes in to the general PIT.
    • LUCI
      • Mostly support for LUCI1 & 2 work and Dave Thompson work in AFC.
      • A note on LUCI and which probaly works with any other instrument: Don't do anything without written confirmation.
      • Handover document
  • Next week:

PK

  • On flight China - Prague: Telemetry C++11
  • populated lib-telemetry
  • OVMS seems to work now
  • Vacation
Next:
  • M1 SX VxWorks replacement
  • 2018B release
Topic revision: r11 - 10 Dec 2018, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback